Salmon often jump waterfalls to reach their
breeding grounds.
Starting downstream, 2.95 m away from a
waterfall 0.54 m in height, at what minimum
speed must a salmon jumping at an angle of
29.7

leave the water to continue upstream?
The acceleration due to gravity is 9.81 m/s
2
.
Answer in units of m/s.
1
Expert's answer
2020-10-01T10:10:49-0400


Remembering that the horizontal component of velocity is constant, the range till the waterfall wall will be


"2.95=v_xt=v\text{ cos}\theta \cdot t."

During the same time of flight, the salmon must reach the height of 0.54 m. This may be before or after the highest point of the trajectory:


"0.54=v_yt-\frac{gt^2}{2}=v\text{ sin}\theta-\frac{gt^2}{2}."

So, we see that we have two equations with two unknowns: "v,t". That is why the system of equations is solvable, solve it with any method you like (the most straightforward is substitution):


"v=7.04\text{ m/s}\\
t=0.483\text{ s}"